OpenStudy (anonymous):

Somehow I think I copied my class notes incorrectly... As part of Rate of Change: At a sand and gravel plant, sand is faling off a conveyor and onto a conical pile at a rate of 10 cubic ft. per min. The diameter of the base of the cone is approximately 3 times the altitude. At what rate is the height of the pile changing when the pile is 15 ft high?

OpenStudy (anonymous):

I don't think I was copying it correctly...I'm confused at what I was doing. \[V = (1/3) \Pi r^2h \] dh/dt = ? when h=15 diameter = 3 * (altitude) 2r = 3h r = (3/2)h \[V = (1/3) \pi ((3/2)h^2)h \] \[V = (3/4)\pi h^3 \] \[dV/dt = (9/4)\pi h^2 (dh/dt)\] \[10 = (9/4) \pi (15)^2 (dh//dt)\] = \[(8/405) \pi ft/\min.\]

OpenStudy (anonymous):

That is a lot. What in particular are you confused about?